Determine whether f : Z × Z → Z is onto if f(m,n)= m-n


1
Expert's answer
2021-12-21T18:27:05-0500

The mapping f : Z × Z → Z , where f(m,n)= m-n

is a surjective function or onto function because for "\forall(z\in Z)" the pair "(z,1)\in dom(f)=Z \times Z" and "f(z,1)=z\cdot 1=z" . So im(f)=Z.